OpenStudy (anonymous):

Solve the equation. Check your answer. 4y+5+-31

OpenStudy (anonymous):

4y+5-31=0?

OpenStudy (anonymous):

4y+5-31=0 4y-26=0 4y=26 y=13/2 y=6.5

OpenStudy (anonymous):

how do you check your answer though?

OpenStudy (anonymous):

Plug what you got for y back into the original equation...

OpenStudy (anonymous):

and y=-9

OpenStudy (anonymous):

can you show me please how to do this one

OpenStudy (anonymous):

4(13/2)+5-31=0 2(13)+5-31=0 26-26=0

OpenStudy (anonymous):

ok put the value of y=13/2 in the equation So now the equation becomes 4*(13/2)+5-31 (4*13)/2=52/2=26 Now 26+5-31=0 31-31=0 0=0 which is LHS=RHS=0

OpenStudy (anonymous):

so y=-9 and that is how you check your answer????

OpenStudy (anonymous):

y is not =-9... y=13/2 When you plug it back into the equation you should get 0

OpenStudy (anonymous):

because it's equal to 0.

OpenStudy (anonymous):

y is not equal to -9 the value of y is 13/2

OpenStudy (anonymous):

Ok tell me is 1= 0 right

OpenStudy (anonymous):

I got y=-9 right on my test though i just needed to check my answer

OpenStudy (anonymous):

y cannot be equal to -9 that is the answer

OpenStudy (anonymous):

worng answer right answer is y=13/2

OpenStudy (anonymous):

so what all should I put on my test then as the whole answer?? because I am so confused

OpenStudy (anonymous):

Did you write the question down right? From your test, are you able to double check it.

OpenStudy (anonymous):

yes i did and I dont understand a bit of it

OpenStudy (anonymous):

According the equation which you have given my answer is the only right answer do you agree @purplec16

OpenStudy (anonymous):

ok so y=13/2 and now can you tell me what is the answer to checking it

OpenStudy (anonymous):

Yes I agree @ipm1988 Maybe she wrote down the question wrong... or your teacher may have given you marks for partial working can you write down your working here...

OpenStudy (anonymous):

no that is the exact question from the test and the qustion again is Solve the equation and check your answer. 4y=5+-31

OpenStudy (anonymous):

opps I mean 4y+5=-31

OpenStudy (anonymous):

-_-, Well actually that's not the exact question... you had 4y+5-31 and you said solve which implies it must be 4y+5-31=0 and not 4y+5=-31... 4y=-36 y=-9

OpenStudy (anonymous):

So to check 4(-9)+5=-31... and you should be able to solve that or plug it into your calculator...

OpenStudy (anonymous):

sorry I had it written right on my paper i just typed it wrong but the real question is solve the equation and check your answer 4y+5=-31

OpenStudy (anonymous):

Well I did that above...

OpenStudy (anonymous):

so the answer is y=-9 and to check it the answer is 4(-9)+5+-31

OpenStudy (anonymous):

your answer -9 is correct here is how you check it but first answer can be 1=0 ?

OpenStudy (anonymous):

no it can't

OpenStudy (anonymous):

Why are you leaving out your equal signs... @jess.white please, be careful

OpenStudy (anonymous):

I mean y=-9 and to check it the answer is 4(-9)+5=-31

OpenStudy (anonymous):

@purplec16 could you answer me back

OpenStudy (anonymous):

no i will not muahahhahaha

OpenStudy (anonymous):

Well, you will not because your not @purplec16... but I stepped away from the computer for a bit, but yes when you solve this equation you will see that they are equal to -31=-31... and that how you know your answer is correct. :).